OpenStudy (anonymous):

Use addition or substitution to find the value of y for this set of equations. 2x - 3y = -15 x = 4y

OpenStudy (anonymous):

your job is to solve \[2\times 4y-3y=-15\]

OpenStudy (anonymous):

8y-3y=-15..

OpenStudy (anonymous):

x = -12, y = -3

OpenStudy (anonymous):

sub x in eqn (1), 5y = -15, y = -3

OpenStudy (anonymous):

x = 4y, x = 4 * -3 = -12